### v3

@@ 2.2 Analysis of Analog Memory Specifications
To verify different analog memory specifications within a neural network system, a weight bank architecture, depicted in Figure 4, is emulated to assess analog memory performance during inference and training using the MNIST dataset [24].
Figure 4: The neural network emulation architecture and initial training. a) The feedforward neural network is trained using the MNIST dataset with 50000, 1-bit, 28x28 pixel images using a batch size of 64, then validated and tested using 10000 images [24]. The neural network is three layers with the input layer supporting 784 values for the 28x28 pixel images, the hidden layer supporting 50 neurons using a ReLU activation function, and the output layer supporting 10 neurons using a logarithmic softmax activation function to classify as a number. The neural network can map to the proposed neuromorphic photonic hardware that consists of 10 photonic cores with 50 rows of weight banks and 80 MRRs in each weight bank. Each core uses 80 wavelengths and 50 semiconductor optical amplifiers (SOA) to compensate for the splitting loss. b) The neural network achieves more than 95% inference accuracy after one epoch for all numbers. The specifics of the neural network architecture, performance, and modeling are detailed in Appendix F.
@@
Figure 5d examines the impact of analog memory retention time constant (weight decay) and network latency on inference accuracy when it degrades by 10%. The weight decay is modeled as an exponential decay based on the ratio of retention time constant to network latency. These experiments compare scenarios where weights are trained with and without considering retention time and network latency. The findings suggest that training with retention time constant and network latency relaxes the requirements for analog memory retention time, enabling the use of memories with shorter retention time. A ratio of retention time constant to network latency greater than 100 prevents inference accuracy from decaying by 10%. This implies that even with an analog memory retention time constant of 100 $\mu$s, if the network latency is less than 1 $\mu$s, the network can maintain high inference accuracy. The ratio of retention time constant to network latency can be scaled to any analog memory technology, giving either a minimum retention time constant for a given network latency or a maximum network latency for a given retention time constant. Although analog memory leakage may limit inference accuracy, controlled leakage can be employed as a regularization technique to avoid overfitting [41]. Furthermore, retention time constant and network latency are varied to evaluate training accuracy with different batch sizes, including 16, 32, and 64. The results show that larger batch sizes require longer retention times to achieve similar training accuracy as smaller batch sizes. With smaller batch sizes, weights are updated more frequently so the analog memory does not need to retain the data for a long time and can operate with shorter retention times.
@@
Once weights become unusable due to leakage, dead-time is consumed to refresh the weight, which can be approximated by the analog memory’s write time. When examining the ratio of retention time to write time for analog memories, this ratio (10,000 for DEOAM) is a couple orders of magnitude greater than the retention time to network latency ratio (100) required to have high inference accuracy. This means the dead-time for refreshing can be much shorter than network latency. Therefore, refreshes can occur after each image classification, thereby maintaining accurate weights for high inference accuracy without adding significant latency but at the cost of refresh power.
@@
Figure 5: Neural network emulation results with and without non-idealities from the analog memory and system, such as control bit precision, noise, leakage, and latency. a) Minimum analog memory control bit precision to achieve more than 95% inference accuracy is 4 bits for weights trimmed only in inference and 8 bits for weights trimmed in inference and training. Inference accuracy degrades with increasing laser, SOA, PD, and TIA noise, but inference accuracy degrades b) earlier when weights are trained without noise and c) later when weights are trained with noise. d) The sweep of network latency and retention time characterizes inference accuracy for weights trained with and without leaky analog memory and network latency. Network latency and analog memory retention time constant at the point when inference accuracy degrades by 10% is plotted. When weights are trained with the effects of network latency and memory retention time, the network compensates for these effects and allows for a lower analog memory retention time constant requirement. Batch sizes of 64, 32, and 16 reveal that large batch sizes require longer retention times since weights are updated less frequently. Weights are updated more frequently in smaller batch sizes. Appendix F expands on results in d).
@@ 3 Discussion
Training neuromorphic photonic processors presents several challenges related to energy efficiency, time, and execution. Analog-to-digital and digital-to-analog conversions along with memory access often become energy and latency bottlenecks. Additionally, since modeling hardware and environmental fluctuations are challenging, relying on offline-trained weights can degrade inference accuracy [42, 11, 9].
To address these challenges, co-locating analog memory with the compute unit can alleviate energy consumption and latency by reducing data movement and the need for ADCs and DACs. A promising outcome from integrating analog memory in neuromorphic photonic processors, as demonstrated in Figure 6, is the potential for fast, efficient, on-chip, online training. This training method uses backpropagation hardware and allows for immediate weight updates with each new input data, reducing the need for analog-to-digital and digital-to-analog conversions. As a result, the processor can adapt swiftly to changing environments and be more resilient to noise and hardware flaws while consuming less energy. Recurrent neural networks, in particular, can benefit from analog memory integration since retaining data in the analog domain reduces analog-to-digital and digital-to-analog conversions [23]. Furthermore, analog memory can enhance the effectiveness of other training algorithms, such as direct-feedback alignment and in-situ training, by improving training time and on-chip computations [43, 11]. Architecturally, incorporating analog memory into neuromorphic photonic processors facilitates system-level improvements, leading to improved energy efficiency, shorter training time, and a more adaptive neural network.
Figure 6: A potential architecture for an fast, efficient neuromorphic photonic processor using analog memory and specialized on-chip, training hardware.
However, scaling the neuromorphic photonic processor with analog memory presents challenges in analog memory and MRR control. Depending on the analog memory requirements, specialized circuitry may be necessary for writing to or refreshing the memory. As the system scales, strategies like sharing drive lines (as illustrated in Figure 1a) where DACs are shared along columns, can be essential in reducing power consumption. Furthermore, as the network grows with more weights, memory refresh rate becomes limited to DAC speeds and the size of the network. Accurate resonance control in MRRs is vital for maintaining precise optical weight representation. As the processor scales, implementing thermal control and multi-MRR control algorithms becomes increasingly critical for system stability [44, 45, 46, 47, 39, 48].
@@
The modulator is another area for enhancement. Employing PN junctions with different doping levels and junction geometries, such as interleaved, U-shaped, or L-shaped designs, can improve tuning efficiency and speed [45, 49, 50]. However, as indicated in Figure 3a, the PN junction MRR exhibits leakage that is dependent on the incident optical power, suggesting that the photoresponsivity of the PN junction can limit analog memory performance. Reducing this leakage may involve removing contaminants and impurities, such as layering poly-silicon on the substrate or optimizing convection annealing [51, 52]. Alternatively, other modulators, such as MOS capacitor modulators, can be explored to potentially reduce leakage [53]. Pockels effect modulators might offer advantages over PN junctions, as they exhibit minimal leakage (less than 100 nA) and better tuning efficiency [54, 55, 56]. Further development in foundry processes is needed to better integrate these modulators into electro-optic platforms [57].
4 Conclusion
@@
We demonstrated a proof-of-concept neuromorphic photonic circuit with DEOAM and verified its capabilities in a neuromorphic system. DEOAM reduces DACs in a neuromorphic photonic processor from n2 to n, thereby alleviating bottlenecks in bandwidth and energy consumption. By integrating analog memory on chip, neuromorphic photonic processors have the potential to achieve fast and efficient training. Since the training process occurs on chip, the neural network can readily adapt to environmental fluctuations and noise by incorporating them during training. Moreover, analog memory with on-chip backpropagation hardware can enable online training, allowing weights to be updated after each new input data. This opens up possibilities for various applications requiring recurrent events, event-based decision-making, or adaptability, such as time series forecasting. As a result, the next generation of AI processors, equipped with integrated analog memory, can help sustain the continued growth of AI.
Acknowledgments

### v5

@@ 2.2  Analysis of Analog Memory Specifications
To verify different analog memory specifications within a neural network system, a weight bank architecture, depicted in Figure 4, is emulated to assess analog memory performance during inference and training using the MNIST dataset [lecun_gradient-based_1998]. The neural network architecture is configured as a three-layer model. The input layer processes 784 values corresponding to the $28\times 28$ pixel image. This is followed by a hidden layer with 50 neurons and a ReLU activation function, and finally, an output layer with 10 neurons and a logarithmic softmax activation for digit classification. The network achieves over 95% validation and testing accuracy across all predicted digits within one training epoch, as shown in Figure 4b, which is comparable to the photonic neural network demonstrated in [ma_high-density_2022]. For a 5-bit resolution MRR-based weight bank matrix incorporating SOAs and MRRs with a finesse of 368 ($\approx$55000 quality factor), the network size is limited to 108 rings (wavelength channels) per weight bank due to the MRRs’ FSR and finesse. The number of rows (spatial channels) is limited to 60 due to the signal-to-noise ratio (SNR) constraints [tait_silicon_2018, al-qadasi_scaling_2022]. To ensure compatibility within a single core, the implementation uses 80 MRRs and 80 analog memories arranged in 50 rows of weight banks. Each weight bank requires two PDs and one TIA to perform the summation and signal amplification.
Figure 4: The neural network emulation architecture and initial training. a) The feedforward neural network is trained using the MNIST dataset with 50000, 1-bit, 28x28 pixel images using a batch size of 64, then validated and tested using 10000 images [lecun_gradient-based_1998]. The neural network is three layers with the input layer supporting 784 values for the 28x28 pixel images, the hidden layer supporting 50 neurons using a ReLU activation function, and the output layer supporting 10 neurons using a logarithmic softmax activation function to classify as a number. The neural network can map to the proposed neuromorphic photonic hardware that consists of 10 photonic cores with 50 rows of weight banks and 80 MRRs in each weight bank. Each core uses 80 wavelengths and 50 semiconductor optical amplifiers (SOA) to compensate for the splitting loss. b) The neural network achieves more than 95% inference accuracy after one epoch for all numbers. The specifics of the neural network architecture, and modeling are detailed in Methods.
@@
Figure 5d examines the impact of analog memory retention time constant (weight decay) and network latency on inference accuracy when it degrades by 10%. The weight decay is modeled as an exponential decay based on the ratio of retention time constant to network latency. These experiments compare scenarios where weights are trained with and without considering retention time and network latency. The findings suggest that training with retention time constant and network latency relaxes the requirements for analog memory retention time, enabling the use of memories with shorter retention time. A ratio of retention time constant to network latency greater than 100 prevents inference accuracy from decaying by 10%. This implies that even with an analog memory retention time constant of 100 $\mu s$, if the network latency is less than 1 $\mu s$, the network can maintain high inference accuracy. The ratio of retention time constant to network latency can be scaled to any analog memory technology, giving either a minimum retention time constant for a given network latency or a maximum network latency for a given retention time constant. Although analog memory leakage may limit inference accuracy, controlled leakage can be employed as a regularization technique to avoid overfitting, enabling broader inference applicability and meta-learning [singhal_enhanced_2023, weilenmann_single_2024]. Furthermore, retention time constant and network latency are varied to evaluate training accuracy with different batch sizes, including 16, 32, and 64. The results show that larger batch sizes require longer retention times to achieve similar training accuracy as smaller batch sizes. With smaller batch sizes, weights are updated more frequently so the analog memory does not need to retain the data for a long time and can operate with shorter retention times.
@@
Once weights become unusable due to leakage, dead-time is consumed to refresh the weight, which consists of the time needed to update the DAC inputs, analog memory write time, and DAC settling time. For DEOAM, the write time is about 65 ns in the fastest case. Surveyed DACs, in similar process nodes to DEOAM, can achieve sample rates up to 3 GS/s (or 333 ps per sample), and reported SRAM propagation delays are in the 10s of picoseconds, meaning that the time needed to update the DAC inputs and DAC settling time is negligible compared to the DEOAM write time [wu_130_2008, kumar_impact_2021]. Even for the fastest write time among surveyed analog memory technologies (500 ps for PCMs), the analog memory write time is similar to DAC sample rates [miller_optical_2018]. Therefore, the dead-time can be approximated by the analog memory’s write time. When examining the ratio of retention time to write time for analog memories, this ratio (10,000 for DEOAM) is a couple orders of magnitude greater than the retention time to network latency ratio (100) required to have high inference accuracy. This means the dead-time for refreshing can be much shorter than network latency. Therefore, refreshes can occur after each image classification, thereby maintaining accurate weights for high inference accuracy without adding significant latency but at the cost of refresh power.
@@
Figure 5: Neural network emulation results with and without non-idealities from the analog memory and system, such as control bit precision, noise, leakage, and latency. a) Minimum analog memory control bit precision to achieve more than 95% inference accuracy is 4 bits for weights trimmed only in inference and 8 bits for weights trimmed in inference and training. Inference accuracy degrades with increasing laser, SOA, PD, and TIA noise, but inference accuracy degrades b) earlier when weights are trained without noise and c) later when weights are trained with noise. d) The sweep of network latency and retention time characterizes inference accuracy for weights trained with and without leaky analog memory and network latency. Network latency and analog memory retention time constant at the point when inference accuracy degrades by 10% is plotted. When weights are trained with the effects of network latency and memory retention time, the network compensates for these effects and allows for a lower analog memory retention time constant requirement. Batch sizes of 64, 32, and 16 reveal that large batch sizes require longer retention times since weights are updated less frequently. Weights are updated more frequently in smaller batch sizes. The Supplementary Information expands on results in d).
@@ 3  Discussion
Training neuromorphic photonic processors presents several challenges related to energy efficiency, time, and execution. Analog-to-digital and digital-to-analog conversions along with memory access often become energy and latency bottlenecks. Additionally, since modeling hardware and environmental fluctuations are challenging, relying on offline-trained weights can degrade inference accuracy [buckley_photonic_2023, filipovich_silicon_2022, sun_full_2023].
To address these challenges, co-locating analog memory with the compute unit can alleviate energy consumption and latency by reducing data movement and the need for ADCs and DACs. Recurrent neural networks, in particular, can benefit from analog memory integration since retaining data in the analog domain reduces analog-to-digital and digital-to-analog conversions [howard_photonic_2020]. Furthermore, analog memory can enhance the effectiveness of other training algorithms, such as direct-feedback alignment and in-situ training, by improving training time and on-chip computations [hughes_training_2018, filipovich_silicon_2022]. A potential outcome from integrating analog memory in neuromorphic photonic processors is on-chip, online training, where the processor can adapt to new data coming in, be more resilient to noise and hardware flaws, and consume less energy [buckley_photonic_2023]. Architecturally, incorporating analog memory into neuromorphic photonic processors facilitates system-level improvements, leading to improved energy efficiency, shorter training time, and a more adaptive neural network.
However, scaling the neuromorphic photonic processor with analog memory presents challenges in analog memory and MRR control. As the network scales, reasonable sizes of the network can be up to 108 rings in a weight bank (limited by the ring’s FSR and finesse) and up to a few hundred rows of weight banks (limited by SNR)[tait_silicon_2018, al-qadasi_scaling_2022]. The size of the network can also be limited to the analog memory’s retention time, refresh time, and inference time (network latency) where a retention time must be allocated to refresh time for all rows of the network and inference time to ensure the network can operate. For DEOAM, the retention time to write time ratio is about 10,000, and inference times (picosecond to nanosecond range) are negligible compared to the write time. Therefore, if SNR is not limiting the number of rows in the network, the next limitation is the refresh rate, which is limited to several thousand rows. Depending on the analog memory requirements, specialized circuitry may be necessary for writing to or refreshing the memory. As the system scales, strategies like sharing drive lines (as illustrated in Figure 1a) where DACs are shared along columns, can be essential in reducing power consumption. Accurate resonance control in MRRs is vital for maintaining precise optical weight representation. As the processor scales, implementing thermal control and multi-MRR control algorithms becomes increasingly critical for system stability [derose_thermal_2014, moazeni_40-gbs_2017, ren_continuously_2018, li_3-d-integrated_2021, tait_feedback_2018, zhang_towards_2014].
@@
The modulator is another area for enhancement. Employing PN junctions with different doping levels and junction geometries, such as interleaved, U-shaped, or L-shaped designs, can improve tuning efficiency and speed [moazeni_40-gbs_2017, liu_silicon_2022, yong_u-shaped_2017]. However, as indicated in Figure 3a, the PN junction MRR exhibits leakage that is dependent on the incident optical power, suggesting that the photoresponsivity of the PN junction can limit analog memory performance. Reducing this leakage may involve removing contaminants and impurities, such as layering poly-silicon on the substrate or optimizing convection annealing [itokawa_reduction_2009, horiuchi_reduction_1995]. Alternatively, other modulators, such as MOS capacitor modulators, can be explored to potentially reduce leakage [hsu_mos_2022]. Pockels effect modulators might offer advantages over PN junctions, as they exhibit minimal leakage (less than 100 nA) and better tuning efficiency [chmielak_pockels_2011, azadeh_measurement_2015, eltes_batio3-based_2019]. Further development in foundry processes is needed to better integrate these modulators into electro-optic platforms [shekhar_roadmapping_2024].
@@
We demonstrated a proof-of-concept neuromorphic photonic circuit with DEOAM and verified its capabilities in a neuromorphic system. DEOAM reduces DACs in a neuromorphic photonic processor from $n^{2}$ to $n$, thereby alleviating bottlenecks in bandwidth and energy consumption. By integrating analog memory on chip, neuromorphic photonic processors have the potential to achieve fast and efficient training. Since the training process occurs on chip, the neural network can readily adapt to environmental fluctuations and noise by incorporating them during training. Moreover, analog memory with on-chip backpropagation hardware can enable online training, allowing weights to be updated after each new input data. This opens up possibilities for various applications requiring recurrent events, event-based decision-making, or adaptability, such as time series forecasting. As a result, the next generation of AI processors, equipped with integrated analog memory, can help sustain the continued growth of AI.
4  Methods
4.1  Hardware
In Figure 2a, the silicon photonic electronic chip, fabricated in the 90 nm GF9WG monolithic process, consists of a 1x4 MRR circuit with C-band grating couplers to couple light to the chip and electronics to control when to load or reset voltages. The electronics consist of transmission gates and reset (RST) transistors. The transmission gates are controlled by an active low enable (EN) signal which connects (EN=0) or disconnects (EN=1) an external source measure unit (SMU) to the analog memory cells. The RST transistors reset the voltage on the analog memory cells before operation. From Figure 2c, the silicon photonic electronic chip is thermally glued and wirebonded to a custom printed circuit board (PCB) that breaks out connections to measurement equipment. The PCB has 50 $\Omega$ transmission lines that breakout to subminiature version A (SMA) connectors for high speed signals and 2.54 mm pitch headers for power and low speed signals. A 10k negative temperature coefficient (NTC) thermistor and peltier module are connected to the LDC501 for thermal control. An 8-degree, 8-channel fiber array is used to couple light to the chip.
4.2  Measurement Setup
From Figure 2a, the optical characteristics are measured using a C-band tunable laser (HP81682A) and an array of photodetectors (N7744A) while the MRRs are voltage biased using a SMU (Keithley 2604B). Light is coupled through an 8-degree, 8-channel fiber array to C-band grating couplers on chip, and the SMU is sweeps voltage to observe MRR resonance changes with varying voltage biases.
For the retention time measurement setup, the HMP4040 power supply unit (PSU) powers the electronics, the LDC501 thermoelectric cooler (TEC) regulates the chip temperature, and the HP81682A tunable C-band laser provides optical power to the circuit for optical processing. To drive signals, the Keithley 2604B SMU is used to drive the analog memory and PN junction MRR and the Agilent 33250A arbitrary waveform generator (AWG) drives the EN signal with a 5 ns edge rate step to quickly connect and disconnect the SMU from the analog memory and PN junction MRR. The DROP port is connected to a PD on the N7744A to tune the laser wavelength to the MRR resonance, and the THRU port is connected to the high speed, amplified PDA255 and DSO81304A oscilloscope to measure the retention time of the circuit. The retention time measurement sequence is shown in Figure 3b. At t0, EN is active, and the SMU sets the desired voltage on the analog memory and PN junction MRR. At t1, EN becomes inactive and the analog memory retains the data on the PN junction MRR. After t1, charge leaks away from the analog memory, which returns the optical response on the THRU and DROP ports back to its original state (back to the MRR’s resonance). At t2, all charge is depleted on the analog memory and the optical response returns to the MRR’s resonance. The leakage measurement setup is the same as the retention time setup, except that the EN signal is constantly active and the SMU is sweeping voltage from 0 V to 2 V while measuring leakage current through the circuit.
The write time measurement setup is similar to the retention time and leakage measurement setup except that the AWG is driving the analog memory and PN junction MRR while the SMU is holding the EN signal active. The write time measurement sequence is shown in Figure 3c and 3e. The AWG sends a pulse train with 5 ns edge rates to drive the analog memory and PN junction MRR, which modulates the optical response on the THRU and DROP port. The optical response is detected using the THRU port by the PD255 and measured by the DSO81304A oscilloscope. The write time measurement sequence shown in Figure 3d differs from Figure 3c and 3e because the rather than driving the analog memory and PN junction with the 5 ns edge rate pulse train, the transmission gate is driven with a pulse train that connects and disconnects the voltage source from the analog memory and PN junction. The rise time, fall time, and one time constant values are extracted from the oscilloscope time domain response.
4.3  Neural Network Architecture
To map the neural network onto hardware, the input data consists of $28\times 28$ 1-bit images (784 pixels), which are mapped to 80 input data modulators per core across 10 cores, for a total of 800 modulators, each handling a single pixel (with some leftover). Each pixel (input data point) is multiplied by 50 different weights to connect to the hidden layer. Since each core contains 50 rows of weight banks, there are 50 MRRs or weights associated with each pixel. Once an input pixel is multiplied by a weight, summation occurs at the PD within each core and then across cores to obtain the 50 neurons in the hidden layer. In each core, 80 pixels are processed (multiplied by their weights) and summed at the PD in each row of weight banks. To process all 784 pixels for the hidden layer, the summation from a given row in one core is combined with the corresponding row in the other nine cores. This yields the total weighted sum for that row, effectively multiplying 50 weights by each of the 784 pixels to produce the inputs to the 50 hidden neurons. A ReLu activation function is assumed to be implemented in analog electronics. The resulting 50 neurons generate 50 intermediate signals, which are passed to the input data modulators for processing in the output layer. In this stage, each signal is multiplied by 10 weights (10 MRRs), allowing the entire multiplication and summation to fit within a single core. After summation, a logarithmic softmax activation (also assumed to be implemented in analog electronics) classifies the signals to predict the corresponding digit.
4.4  Noise
Laser noise represents input data noise and can be related to relative intensity noise (RIN) from the laser as shown in the following equations [hui_chapter_2023]:
$x_{i}=\overline{x_{i}}(1+dx_{i})=\overline{P_{i}}(1+dP_{i})=\overline{P_{i}}(1+\mathcal{N}(0,10^{RIN/10}\cdot f))$ (3)
where $x_{i}$ is the $i^{th}$ input data, $\overline{x_{i}}$ is the average value of the $i^{th}$ input data, $dx_{i}$ is the $i^{th}$ random noise for the $i^{th}$ input data, $\overline{P_{i}}$ is the $i^{th}$ average laser power, $dP_{i}$ is the $i^{th}$ random noise in the laser power, $\mathcal{N}$ is the normal distribution function, $RIN$ is the $i^{th}$ relative intensity noise in dB/Hz, and $f$ is the integration bandwidth in Hz. As laser light splits to reach each weight bank, splitting loss is incurred. SOAs can amplify the signal to compensate for splitting loss but adds noise to the input data. As a result, the input data to each weight can be described by:
$x_{i}^{\prime}=x_{i}+n_{\mathrm{SOA_{j}}}$ (4)
where $x_{i}^{\prime}$ is the $i^{th}$ input data to the $i^{th}$ weight that includes SOA and laser noise, and $n_{\mathrm{SOA_{j}}}$ is the $j^{\mathrm{th}}$ row SOA noise. As the input data is multiplied by weights, insertion loss from MRRs degrades the input data signal. Weight bank insertion loss ($\beta$) is modeled by using an insertion loss of 0.125 dB/MRR, giving a total insertion loss of 10 dB in each weight bank. At the input of each weight bank, each wavelength has 1 mW of optical power and the noise ranges from 0% to 37.5% of 1 mW to mimic noise from the laser and SOAs. Furthermore, noise is applied to the data after the MAC operation and before the activation function to mimic the input referred noise current on the PD and TIA. The PD and TIA integrated noise current ranges from 1 nA to 2 mA based on current PD and TIA technology [di_patrizio_stanchieri_18_2022, chuah_design_2015, wang_multi-channel_2020, atef_optical_2013, babar_83-ghz_2022, li_112_2018, awny_40_2015, kudszus_46-ghz_2003, awny_235_2016]. The following equation summarizes the noise sources and the neural network equation:
$y_{j}=\sigma_{j}(\beta\mathbf{x^{\prime}}\cdot\mathbf{w}_{j}+n_{j})$ (5)
where $y_{j}$ is the $j^{\mathrm{th}}$ output data in a layer, $\beta$ is the weight bank insertion loss, $\sigma_{j}$ is the $j^{\mathrm{th}}$ activation function in a layer, $\mathbf{x^{\prime}}$ is the input data vector to the weight banks that includes SOA and laser noise ($\mathbf{x^{\prime}}=<x_{0}^{\prime},x_{1}^{\prime},...,x_{i}^{\prime}>$), $\mathbf{w}_{j}$ is the $j^{\mathrm{th}}$ vector of weights in a layer, and $n_{j}$ is the $j^{\mathrm{th}}$ input referred noise term. The dot product of $\mathbf{x^{\prime}}$ and $\mathbf{w}_{j}$ represents the photocurrent from the PD detecting laser wavelengths weighted by MRRs.
4.5  Weight Decay and Network Latency
For each weight, weight decay is modeled by using a ratio of retention time to network latency in an exponential decay model to determine inference accuracy from leaky weights:
$w(k)=w_{0}e^{-k\frac{t_{\mathrm{latency}}}{\tau_{\mathrm{ret}}}}$ (6)
where $w(k)$ is the weight value for the $k^{\mathrm{th}}$ image, $w_{0}$ is the initial weight, $t_{\mathrm{latency}}$ is the network latency or compute time, $\tau_{\mathrm{ret}}$ is the retention time constant. In inference, network latency is determined by group delay from photons propagating through the optical circuit and driver, modulator, PD and TIA bandwidth, whereas training has additional latency from analog electronics used to update weights. Since the bandwidth of the PD and TIA and analog electronics can vary depending on design and technology nodes, the range of network latencies are assumed to be determined by the PD and TIA bandwidth, which translates to latencies ranging from 50 ps to 10 ns [di_patrizio_stanchieri_18_2022, chuah_design_2015, wang_multi-channel_2020, atef_optical_2013, babar_83-ghz_2022, awny_40_2015, kudszus_46-ghz_2003, awny_235_2016].
[…]